### v1

# A Classification-Driven Likelihood Ratio Method for Familial DNA Testing

Abstract
Background Familial DNA search relies on statistical methods to infer genetic relationships, often using the likelihood ratio (LR) as the test statistic. This approach involves calculating the probability of observed DNA profiles under various scenarios, traditionally assuming a uniform genetic background for all individuals. However, this assumption is unrealistic in practice due to population substructure, where allele frequencies differ among subpopulations. To address this, modified LR-based approaches have been proposed, including methods such as calculating likelihood ratios using average allele frequencies (LRLAF) or the maximum, minimum, and average likelihood ratios (LRMAX, LRMIN, and LRAVG). Despite these developments, further advancements are needed to improve the testing of genetic relationships in the presence of subpopulation structures.
Results We propose a novel LR-based statistic designed to account for population substructure by incorporating a classification step to address nuisance parameters. Specifically, two DNA profiles with unknown subpopulation origins are simultaneously classified into one group. Following this classification, the likelihood ratio (LRCLASS) is applied. Our analysis demonstrates that the proposed LRCLASS statistic, when paired with Naive Bayes classification, exhibits higher statistical power than existing methods for testing full-sibling relationships in the Thai population. Additionally, we explored an alternative classification approach using multinomial logistic regression, which achieved performance comparable to the Naive Bayes method.
Conclusion The proposed LRCLASS statistic enhances the power of familial DNA searches by effectively addressing population substructure through a novel classification-based approach. This method provides a robust alternative to existing LR-based strategies and underscores the potential for integrating advanced classification techniques into genetic relationship testing. Future research could investigate other classification methods in conjunction with likelihood ratio calculations to further improve the accuracy and applicability of familial DNA analyses.
keywords:
familial DNA search, likelihood ratio, population substructure, classification
1 Background
1.1 Familial DNA Search
In forensic investigations, genetic material or DNA is one of the most common types of evidence collected from crime scenes. It can be extracted from various parts of the human body, such as hair, blood, sweat, or saliva. A typical question in such cases is whether a DNA sample from the crime scene matches any individual in a forensic database. If a match is found, the investigation becomes more straightforward, as the individual who left the sample can be identified. The results can then be submitted to the court for further proceedings. However, if the DNA sample from the crime scene does not perfectly match anyone in the database but shows some similarities, it may indicate a genetic relationship between two individuals, such as parent-child or full-sibling. This situation leads to familial DNA searches, which use statistical tools to infer familial relationships based on DNA evidence. Various strategies for familial searching have been proposed in the literature, ranging from basic methods that count the number of shared alleles to more advanced probabilistic approaches (e.g., Balding et al. 2013, Bieber et al. 2006, Ge et al. 2011, Ge & Budowle 2012, Kooakachai et al. 2019, Kruijver et al. 2014, Slooten & Meester 2014). Familial DNA searches have grown in prominence over the years. By 2014, at least 40 laboratories in the United States were conducting such searches, and approximately 75% of those not yet practicing this method were actively considering its implementation (Debus-Sherrill & Field, 2019). This approach is one of several indirect methods for identifying potential sources of forensic biological evidence, alongside techniques like Y-STR database searches and investigative genetic genealogy, as discussed by Ge and Budowle (2021). Despite its utility, familial DNA searching has sparked ethical concerns, particularly regarding the disproportionate representation of nonwhite individuals in databases and the potential for intrusive investigations targeting relatives of suspects. In response, U.S. states have adopted varied policies: some have formalized rules permitting or banning familial searches, while others allow follow-up investigations on accidental partial matches under regulatory oversight to mitigate ethical issues (Katsanis 2020).
In statistical hypothesis testing for familial DNA searches, the likelihood ratio statistic is frequently employed to evaluate the probability of a pair of DNA profiles under two distinct relationship hypotheses. One scenario assumes the individuals are unrelated, while the other considers a specific familial relationship, such as parent-child or full-sibling. Over the years, the adoption of the likelihood ratio statistic for relatedness testing has grown significantly. A study on the New Zealand database, for instance, demonstrated that the likelihood ratio approach for parent-child and full-sibling tests outperformed the traditional method of counting matching alleles (Curran & Buckleton 2008). Building on this, a validation study further highlighted the effectiveness of a likelihood ratio-based method for identifying first-degree familial relationships using autosomal STR profiles in California’s 1,000,000-profile database (Myers et al. 2011). This approach was subsequently implemented in California’s Familial Search Program, where it played a pivotal role in solving the Grim Sleeper serial murder case. The likelihood ratio statistic is both straightforward and statistically robust, but it comes with notable limitations. A key drawback is its assumption that all individuals in the population share the same genetic background, meaning their genetic information is drawn from a single allele frequency distribution. However, this assumption is often unrealistic in forensic applications. Populations are more accurately modeled as having multiple genetic substructures, such as racial or ancestral groups, which can significantly influence allele frequencies. Failing to account for these differences can lead to inaccurate likelihood ratio calculations and erroneous conclusions, potentially resulting in miscarriages of justice (Balding & Nichols 1994). In this context, population substructure refers to the genetic variation that exists between distinct subgroups within a larger population. These variations must be incorporated into likelihood ratio calculations to ensure accuracy and fairness in forensic analyses.
The critical role of subpopulation structure in forensic science is emphasized in numerous studies to ensure accurate DNA analysis and minimize errors. Factors such as geographical proximity, migration patterns, and cultural practices are shown to shape subpopulation structures, significantly influencing likelihood assessments in forensic investigations (Gallo et al. 1997). Allelic frequencies across diverse racial and ethnic databases have been examined, revealing substantial variations in probability calculations (Gallo et al. 1997). Differences in allele frequencies among ethnic groups, combined with genetic admixture, have been demonstrated to increase the risk of false positives or reduce the statistical power of genetic studies (Tian et al. 2008). Strategies to address population substructure in DNA database searches have been proposed, with simulations using ten-locus DNA profiles showing that even mild substructure can inflate evidence strength in 11.6% of cases (Spooner & Stockmarr 2019). A comprehensive overview of the importance of subpopulation structure has been provided, including discussions on foundational statistical concepts and recent advances, such as methods for ancestral classification using informative single nucleotide polymorphism markers (Jeong et al. 2021).
When allele frequencies are available for each subpopulation, adjustments to the likelihood ratio formula have been proposed by various researchers. One notable approach, implemented by the Denver Crime Laboratory in Colorado, USA, in 2008, was inspired by the work of Prescutti et al. (2002). This method involves calculating likelihood ratios multiple times, applying a single population framework to each available allele frequency distribution. The highest value from these calculations, referred to as LRMAX, is then selected as the final likelihood ratio statistic. Alternatively, a different approach, introduced by Ge and Budowle (2012), suggests using the minimum likelihood ratio instead. This method argues that the minimum likelihood ratio provides a more accurate kinship index for the population associated with the forensic profile. Their research demonstrated that the minimum likelihood ratio consistently outperformed the maximum method in terms of accuracy. Two additional likelihood ratio-based statistics were introduced in 2019. The first, LRLAF, applies the classical likelihood ratio framework using average available allele frequencies, while the second, LRAVG, calculates the average of all likelihood ratios derived from each allele frequency distribution (Kooakachai et al. 2019). When these methods were compared for parent-child and full-sibling tests in the Colorado population, LRLAF was found to exhibit approximately 10% higher power for parent-child tests and 5% higher power for full-sibling tests compared to LRMAX. However, debate continues regarding the consistency of these methods, particularly when applied to populations outside of Colorado.
Existing methods for familial DNA searches in the literature primarily rely on basic arithmetic, without incorporating advanced classification techniques. Instead of simply averaging likelihood ratios (LRAVG), averaging at the allelic level (LRLAF), or using the maximum (LRMAX) or minimum (LRMIN), machine learning methods can be employed to predict the subpopulation to which a DNA profile belongs, followed by a likelihood ratio calculation after classification. The use of classification has already been applied in some forensic subfields. For example, the application of penalized logistic regression in forensic toxicology was demonstrated for calculating likelihood ratios in a two-class classification setting (Biosa et al. 2020). The effectiveness of this framework was illustrated through a case study on alcohol biomarker data, highlighting its ability to distinguish chronic alcohol drinkers and its adaptability for use in forensic and analytical chemistry applications.
This study introduces a novel likelihood ratio-based statistic designed to assess the genetic relationship between two individuals. The proposed method leverages a Naive Bayes classification framework, wherein DNA samples are initially categorized into subpopulations, and the likelihood ratio is subsequently computed based on the resulting classifications. The efficacy of the proposed statistic will be rigorously evaluated through its Type I error rates and statistical power. Furthermore, an alternative classification approach employing multinomial logistic regression is introduced, enabling a comparative analysis of the classification performance between the Naive Bayes and multinomial logistic regression methodologies.
1.2 DNA Profile and Dataset Description
Humans differ due to the random combination of nucleobases—adenine (A), thymine (T), guanine (G), and cytosine (C)—inherited from both parents. The human genome, composed of about 3 billion nucleobases and approximately 20,000 genes across 23 pairs of chromosomes, contains all the genetic instructions for an individual. Genetic markers are variations in DNA sequences at specific positions on chromosomes. Among these markers, short tandem repeats (STRs) are especially useful in forensic science. STRs consist of DNA sequences, typically two to seven base pairs long, repeated multiple times. The number of repeats at each locus varies between individuals, allowing for identification. For example, a 28-base pair sequence like TCTATCTATCTATCTATCTATCTATCTA represents seven repeats of the four-base sequence TCTA, indicating an allele of length 7 at that position. Each person inherits one STR allele from each biological parent, and the number of repeats can differ between them. A genotype at a specific locus is represented by two alleles, such as 11,12, meaning the individual has 11 and 12 repeats of the TCTA sequence at that locus. The variation in STR repeat lengths among individuals makes STRs highly effective for forensic identification.
DNA databases, including government repositories, play a crucial role in criminal investigations by helping identify potential suspects. Since their inception, forensic databases have largely depended on STR markers due to their higher variability and stronger discriminatory ability compared to other genetic markers, which makes them vital for precise identification (Butler 2005).
In this paper, a DNA profile is described as the collection of genotypes across multiple loci for an individual, with the number of loci used in forensic investigations differing by region. Here, Thailand is highlighted as a case study due to its recognized genetic substructure and documented genetic diversity, as supported by archaeological and linguistic research. Genetic differences have been observed between Thai-Malay Muslims, primarily in southern Thailand, and Thai-Buddhists from other regions (Kutanan et al. 2014a). Distinct genetic profiles have also been identified between Thai-northeastern populations and other groups, including those in northern Thailand, through mitochondrial DNA-HVR1 variation studies conducted in the same year (Kutanan et al. 2014b). More recently, significant genetic differences among subpopulations have been emphasized through a genome-wide analysis of 32 Thai subpopulations, underscoring the importance of careful participant selection for biomedical and clinical research (Kutanan et al. 2021). Although Thailand does not yet have a National DNA Database, a DNA database was established by the Central Institute of Forensic Science (CIFS) in 2004 to address crimes in the Deep South. This database, which has since expanded to include approximately 160,000 profiles of crime samples, individuals, and prisoners, utilizes autosomal and Y-STR markers. A 24-loci STR kit was introduced in 2016 to support both direct and familial matching (Boonderm et al. 2017).
To simplify the analysis, four main Thai subpopulations, defined by geographical regions—North (NO), Northeastern (NE), Central (CT), and South (SO)—are considered in this study. Allele frequency distributions for these groups were derived from DNA samples of 929 unrelated Thai individuals analyzed across 15 loci (Shotivaranon et al. 2009). The sample sizes for the North, Northeastern, Central, and South subpopulations were 202, 304, 212, and 211, respectively. Based on the availability of data, 15 core loci are utilized in the study: D8S1179, D21S11, D7S820, CSF1P0, D3S1358, TH01, D13S317, D16S539, D2S1338, D19S433, vWA, TPOX, D18S51, D5S818, and FGA.
2 Methods
The task of determining whether two DNA profiles $X_{1}$ and $X_{2}$ share a specific genetic relationship can be framed as a hypothesis test:
$H_{0}:$ The two individuals are unrelated.
$H_{1}:$ The two individuals share a specific genetic relationship.
The genetic relatedness between two individuals can be described using the notation proposed by Lynch and Walsh (1998). They suggested that all familial relationships can be explained through the probabilities of alleles at a locus being identical by descent (IBD), where two alleles are considered IBD if they are inherited as copies of the same ancestral allele. For any two individuals, the relatedness at a specific locus is expressed as a 3-tuple $\theta$ = $(z_{0},z_{1},z_{2}).$ Here, $z_{0}$ represents the probability that neither allele of $X_{1}$ is IBD to any allele of $X_{2}$, $z_{1}$ refers to the probability that exactly one allele of $X_{1}$ is IBD to an allele of $X_{2},$ and $z_{2}$ indicates the probability that both alleles of $X_{1}$ are IBD to those of $X_{2}$. Using this definition, the parent-child relationship is represented by $\theta$ = (0,1,0), as the parent and child always share exactly one ancestral allele at each locus. Similarly, the unrelated and full-sibling relationships are expressed as $\theta$ = (1,0,0) and (1/4, 1/2, 1/4), respectively. It is important to note that, by definition, the probabilities $z_{0},z_{1},$ and $z_{2}$ must sum to one.
The hypotheses for testing a parent-child relationship can be rewritten as the following:
$H_{0}$ : $\theta$ = (1,0,0)    versus    $H_{1}$ : $\theta$ = (0,1,0),
whereas testing for a full-sibling relationship corresponds to
$H_{0}$ : $\theta$ = (1,0,0)    versus    $H_{1}$ : $\theta$ = ($\frac{1}{4},\frac{1}{2},\frac{1}{4}$).
The test statistic used is the likelihood ratio $\Lambda$ defined by
$\Lambda$ $=\dfrac{P(X_{1},X_{2}\mid H_{1})}{P(X_{1},X_{2}\mid H_{0})}$
If $\Lambda>1$, then there is more chance that the two DNA profiles $X_{1}$ and $X_{2}$ were generated from $H_{1}$ compared to $H_{0}$. In particular, higher $\Lambda$ corresponds to more evidence to reject the null hypothesis and conclude a specific relationship of interest.
Calculating the probabilities in the numerator and denominator of $\Lambda$ requires knowledge of the allele frequency distributions. Let $f_{X_{1}}$ and $f_{X_{2}}$ represent the allele frequency distributions for the populations to which $X_{1}$ and $X_{2}$ belong, respectively. These distributions provide the frequencies of each allele, allowing the probabilities to be fully determined. That is, $f_{X_{1}}$ and $f_{X_{2}}$ are essential inputs for calculating the likelihood ratio, serving as nuisance parameters. Additionally, the probabilities are calculated separately for each locus and then combined by multiplying them, based on the assumption that loci are independent. The assumption of independence between loci requires that the loci are not genetically linked and are in linkage equilibrium. However, some core loci do not fully satisfy this assumption. For instance, the loci vWA and D12S391 are located on the same chromosome, approximately 6.3 megabases (Mb) apart. Ideally, genetic markers on the same chromosome should be separated by more than 50 Mb to ensure independence between genotypes at the two loci (O’Connor et al. 2011). Studies have indicated that independence between vWA and D12S391 cannot always be assumed at the individual level. Nevertheless, at the population level, there is no significant evidence of correlation between genotypes at the core loci, including these two loci (Budowle et al. 2011, Ge et al. 2012). For simplicity in our calculations, we proceed under the assumption of independence between loci.
In the simplest case of a population with no substructure, $f_{X_{1}}$ and $f_{X_{2}}$ are equal and correspond to the allele frequency distribution of the entire population, making the likelihood ratio test well-defined. However, when the population exhibits substructure, $f_{X_{1}}$ and $f_{X_{2}}$ can take on multiple possible values. Currently, there is no established theory identifying the optimal test statistic for such cases. This poses a significant challenge in determining the most powerful statistic for hypothesis testing in populations with substructure.
2.1 Existing Test Statistics
Consider a population divided into $R>1$ distinct subpopulations, each with its own allele frequency distribution, denoted as $f_{i}$ for the $i^{\text{th}}$ subpopulation. Below, we define four existing methods for conducting familial DNA testing when the profiles $X_{1}$ and $X_{2}$ are given.
The first statistic is obtained by using the likelihood ratio framework but plugging in the value of $f_{X_{1}}$ and $f_{X_{2}}$ by $f_{local}$ which is defined by
$f_{local}=p_{1}f_{1}+p_{2}f_{2}+\cdots+p_{R}f_{R}$
where $p_{i}$ is the proportion of individuals in the $i^{\text{th}}$ subpopulation. This statistic is called LRLAF and can be written mathematically as
$\Lambda_{LAF}=\frac{L(\theta_{1}\mid X_{1},X_{2},f_{local})}{L(\theta_{0}\mid X_{1},X_{2},f_{local})}.$
Examining the LRLAF approach reveals that it involves averaging directly at the allele frequency level. A slight modification can be made by shifting the averaging process to the likelihood ratio level instead. This leads to a new statistic called LRAVG. In this method, the likelihood ratios are computed repeatedly using the single-population framework applied to multiple locally available allele frequency distributions. The final step involves calculating a weighted average of these likelihood ratios. Mathematically, the LRAVG statistic is expressed as follows:
$\Lambda_{AVG}=p_{1}\frac{L(\theta_{1}\mid X_{1},X_{2},f_{1})}{L(\theta_{0}\mid X_{1},X_{2},f_{1})}+p_{2}\frac{L(\theta_{1}\mid X_{1},X_{2},f_{2})}{L(\theta_{0}\mid X_{1},X_{2},f_{2})}+\dots+p_{R}\frac{L(\theta_{1}\mid X_{1},X_{2},f_{R})}{L(\theta_{0}\mid X_{1},X_{2},f_{R})}.$
If instead of calculating a weighted average in the final step, we opt to take the maximum or minimum, the resulting statistics are referred to as LRMAX or LRMIN, respectively. The definitions of these two statistics are as follows:
$\Lambda_{MAX}=\max\Bigl{\{}\frac{L(\theta_{1}\mid X_{1},X_{2},f_{1})}{L(\theta_{0}\mid X_{1},X_{2},f_{1})},\frac{L(\theta_{1}\mid X_{1},X_{2},f_{2})}{L(\theta_{0}\mid X_{1},X_{2},f_{2})},\dots,\frac{L(\theta_{1}\mid X_{1},X_{2},f_{R})}{L(\theta_{0}\mid X_{1},X_{2},f_{R})}\Bigr{\}}.$
$\Lambda_{MIN}=\min\Bigl{\{}\frac{L(\theta_{1}\mid X_{1},X_{2},f_{1})}{L(\theta_{0}\mid X_{1},X_{2},f_{1})},\frac{L(\theta_{1}\mid X_{1},X_{2},f_{2})}{L(\theta_{0}\mid X_{1},X_{2},f_{2})},\dots,\frac{L(\theta_{1}\mid X_{1},X_{2},f_{R})}{L(\theta_{0}\mid X_{1},X_{2},f_{R})}\Bigr{\}}.$
2.2 Proposed Test Statistic
The key to successfully computing the likelihood ratio lies in eliminating the nuisance parameters $f_{X_{1}}$ and $f_{X_{2}}$. Existing methods typically rely on basic statistics like the average, maximum, and minimum. However, more advanced statistical techniques can be employed to predict the true values of $f_{X_{1}}$ and $f_{X_{2}}$. An example of this can be seen in the application of Naive Bayes classification.
To understand how Naive Bayes classification operates, consider a population divided into $R>1$ distinct subpopulations. Given a DNA profile $B$, the task is to determine which subpopulation it belongs to. This can be achieved by calculating the probability that the profile $B$ originates from each allele frequency distribution. Mathematically, let $A_{i}$ be the event such that a DNA profile $B$ corresponds to the $i^{\text{th}}$ subpopulation. By using the Bayes’ rule, for i{1, 2,,R},i\ $\in\{$1,\ 2,\ $\ldots$,\ R\}, we have
$P\left(A_{i}\mid B\right)=\frac{P\left(B\mid A_{i}\right)P\left(A_{i}\right)}{\sum_{k=1}^{R}{P\left(B\mid A_{k}\right)P\left(A_{k}\right)}}\ .$
On the right-hand side of the equation, the subpopulation prior probability $P\left(A_{i}\right)$ can be determined using the proportion of each subpopulation. The term $P\left(B\mid A_{i}\right),$ which represents the probability of observing a DNA profile $B$ given that the associated subpopulation is $i,$ can be computed using allele frequency data from the $i^{\text{th}}$ subpopulation. An illustrative example is presented here: consider a scenario with $R=4$ subpopulations and three independent loci, namely TPOX, vWA, and D13S317. The DNA profiles of two individuals, $X_{1}$ and $X_{2}$, are presented in Table 1:
Table 1: An example of DNA profiles
Locus
DNA profiles TPOX vWA D13S317
$X_{1}$ $(10,10)$ $(15,17)$ $(9,10)$
$X_{2}$ $(10,11)$ $(15,15)$ $(9,10)$
[…]
Additionally, assume that the allele frequency distributions for the specified loci are provided in Table 2. Assume further that the prior probabilities for each subpopulation are $P(A_{1})=0.1,P(A_{2})=0.2,P(A_{3})=0.3$ and $P(A_{4})=0.4.$
Table 2: The allele frequency distributions for the specified loci with respect to the example.
Allele Locus prob. subpop1 prob. subpop2 prob. subpop3 prob. subpop4
10 TPOX 0.0223 0.0378 0.0354 0.0569
11 TPOX 0.2351 0.2072 0.2618 0.2867
12 TPOX 0.0173 0.0263 0.0071 0.0237
15 vWA 0.0396 0.0263 0.0283 0.0332
16 vWA 0.1015 0.0872 0.1297 0.1445
17 vWA 0.2475 0.2484 0.2524 0.2725
9 D13S317 0.1386 0.1464 0.1344 0.1398
10 D13S317 0.1609 0.1168 0.1179 0.0995
11 D13S317 0.2327 0.2072 0.2406 0.2512
[…]
We consider the simultaneous classification of $X_{1}$ and $X_{2}$. Given the independence of the two profiles, we have:
$P\left(X_{1},X_{2}\mid A_{i}\right)=P\left(X_{1}\mid A_{i}\right)P\left(X_{2}\mid A_{i}\right).$
To calculate the genotype probability across $m$ loci, we assume that these loci are independent. Therefore, the genotype probability across $m$ loci is the product of the genotype probabilities at each locus. Specifically, for $k=1,2$, we have
$P\left(X_{k}\mid A_{i}\right)=\prod_{j=1}^{m}{P\left(X_{k,j}\mid A_{i}\right)}$
where $X_{k,j}$ represents the genotype of $X_{k}$ at locus $j$. It is important to note that the genotype at locus $j$ can be of two types: homozygous (aa) or heterozygous (ab). The corresponding probabilities for these genotypes are $p^{2}$ and $2pq,$ where $p$ and $q$ are the allele frequencies of alleles a and b at locus $j$, respectively. In the context of this example, we set $m=3$, and an example of the complete calculation of genotype probabilities for the case where $k=1$ and $i=1$ is shown as follows:
$P\left(X_{1}\mid A_{1}\right)$ $=\prod_{j=1}^{3}{P\left(X_{1,j}\mid A_{1}\right)}=(0.0223^{2})(2(0.0396)(0.2475))(2(0.1386)(0.1609))$
$=4.3477\times 10^{-7}$
The calculations for other values of $k$ and $i$ can be performed in the same manner. When all values are combined, the results in this context are $P\left(A_{1}\mid X_{1},X_{2}\right)=0.2903,P\left(A_{2}\mid X_{1},X_{2}\right)=0.0431,P\left(A_{3}\mid X_{1},X_{2}\right)=0.0729,$ and $P\left(A_{4}\mid X_{1},X_{2}\right)=0.5938$.
[…]
$\arg\max_{i}\{P\left(A_{i}\mid B\right)\}.$
Therefore, in this specific example, DNA profile $X_{1}$ and $X_{2}$ are classified into the $4^{\text{th}}$ subpopulation, as it has the highest probability.
The following provides the formal definition of the proposed statistic. Given DNA profiles for two individuals, we aim to simultaneously classify which subpopulation these profiles belong to. In the first step, Naive Bayes classification is applied. Then, the likelihood ratio statistic is used within a single population framework, based on the classification result. This statistic, denoted as LRCLASS, is mathematically expressed as:
$\Lambda_{CLASS}:=\frac{L(\theta_{1}\mid X_{1},X_{2},f_{result})}{L(\theta_{0}\mid X_{1},X_{2},f_{result})}$
where $f_{result}$ represents the subpopulation predicted from the classification.
2.3 Algorithms for Power Comparison
To compare the existing statistics with the proposed statistic, we evaluate them based on statistical power. The power of each statistic will be determined through simulations. These simulations will involve a sample size of 1,000,000 pairs under the null hypothesis and 1,000,000 pairs under the alternative hypothesis. The numerical analysis will be conducted using the Thai population as a case study, focusing on parent-child and full-sibling relationships.
Recall that power refers to the probability of correctly rejecting the null hypothesis when the alternative hypothesis is true. To determine this, we first generate a null distribution to establish a criterion for rejecting the hypothesis, and then create the alternative hypothesis to estimate the power. The algorithm is outlined below.
  1. 1.
    Simulate a pair of unrelated DNA profiles using $m=15$ loci. The subpopulation information of the individuals is randomly selected from a multinomial distribution with success probabilities $p_{NO}=0.11083,\,p_{NE}=0.36944,\,p_{CT}=0.35383,\,p_{SO}=0.16590$. These numbers are estimated from Thai population database (2023).
  2. 2.
    Calculate the likelihood ratio statistics for the pair drawn in Step 1 using five different formulas: LRLAF, LRAVG, LRMIN, LRMAX, and LRCLASS. Save the resulting values.
  3. 3.
    Repeat Steps 1 and 2 many times, for example, $B=$ 1,000,000 times. This will generate the null distribution of likelihood ratios.
  4. 4.
    Estimate a threshold $c$ from the null distribution obtained in Step 3. Specifically, we determine $c$ such that $P(\text{LR}>c)=\alpha$, where $\alpha$ is the specified false positive rate.
  5. 5.
    Generate a pair of DNA profiles based on the relationship in the alternative hypothesis (using $m=15$ loci). For each pair, the subpopulation information of the first individual is selected randomly from a multinomial distribution, as introduced in Step 1, while the subpopulation information of the second individual is forced to match the first, as they are genealogically related.
  6. 6.
    Calculate the likelihood ratio statistics for the pair drawn in Step 5 using five different formulas: LRLAF, LRAVG, LRMIN, LRMAX, and LRCLASS. Save the resulting values.
  7. 7.
    Repeat Steps 5 and 6 multiple times (for instance, 1,000,000 times) to obtain the alternative distribution of likelihood ratios based on the generated samples.
  8. 8.
    Estimate the power by calculating the ratio of the number of pairs with statistics greater than the threshold $c$ to the total number of simulated pairs in Step 7.
Once the power for each statistic is obtained, we aim to compare them at the same false positive rate (FPR) $\alpha$. To determine an appropriate value for $\alpha$, we can refer to previous studies in the Colorado prison population, which suggested that for parent-child and full-sibling tests, $\alpha$ should be $1.7\times 10^{-5}$ and $1.2\times 10^{-5},$ respectively (Kooakachai et al. 2019).
2.4 Classification Comparison
We have developed the LRCLASS statistic, which classifies two DNA profiles into a single subpopulation simultaneously. The classification method employed was the Naive Bayes classifier. However, in statistics and machine learning, numerous other classification techniques are available. In this section, we describe an alternative classification method—logistic regression—and discuss how its performance compares to the previously proposed Naive Bayes classifier.
To begin, two approaches are outlined for representing DNA profiles as explanatory variables in logistic regression. A DNA profile with $m$ loci can be expressed in the general format:
$\{(x_{1a},x_{1b}),(x_{2a},x_{2b}),(x_{3a},x_{3b}),\ldots,(x_{ma},x_{mb})\}$
where $x_{ij}$ denotes the allele observed at locus $i\in\{1,2,\ldots,m\}$ and $j\in\{a,b\}$ is the maternal or paternal origin of the allele. The target variable in the proposed logistic regression is the subpopulation to which an individual belongs. Given the consideration of multiple subpopulations, a multinomial logistic regression model is utilized. The predictors can be structured in two distinct ways:
Method A Treat each allele count $x_{ij}$ as a categorical variable and use them directly as predictors in the regression model. This results in $2m$ categorical variables.
Method B Represent the genotype at each locus as a single categorical variable. Specifically, define $u_{i}=(x_{ia},x_{ib})$ as the predictor for locus $i$. This approach reduces the number of predictors to $m$, but each variable includes a larger number of categorical levels compared to Method A.
Logistic regression will yield the probabilities of each DNA profile belonging to each subpopulation. The predicted subpopulation is then determined as the one with the highest probability.
To evaluate the performance of three classification methods—logistic regression with Method A, logistic regression with Method B, and Naive Bayes classification—we employ the following algorithm:
  1. 1.
    A total of 100,000 Thai DNA profiles are generated based on allele frequency distributions reported by Shotivaranon et al. (2009).
  2. 2.
    The dataset from Step 1 is divided into five equal parts. For each trial, four parts (80% of the dataset) are used as the training set, while the remaining part (20%) is designated as the validation set. This process is repeated five times, ensuring that each part serves as the validation set exactly once. This approach is analogous to the concept of 5-fold cross-validation.
  3. 3.
    In each trial, subpopulation prediction is performed on the validation set using three distinct methods: logistic regression with Method A (requiring the training set), logistic regression with Method B (requiring the training set), and Naive Bayes classification (not requiring the training set). The corresponding validation errors are recorded in the form of a confusion matrix and summarized using standard metrics such as precision, recall, specificity, accuracy, error, and F1-score, which will be further discussed in the subsequent section.
  4. 4.
    The analyses in Step 3 are conducted five times, each based on a different validation dataset, and the average of these five repetitions will be used for comparison.
  5. 5.
    We also utilize the test dataset from Shotivaranon et al. (2009), which includes 929 samples from Thai individuals. Subpopulation prediction is performed on this test dataset using three different methods: logistic regression with Methods A and B (both requiring the 100,000 generated samples from Step 1 as the training data), and Naive Bayes classification (which does not require a training set). The corresponding test errors are then recorded.
  6. 6.
    The performance of the three classification methods is compared by examining the test errors obtained in Step 5.
3 Results
3.1 Power Comparison
Table 3 displays the critical thresholds for rejecting the null hypothesis associated with each test statistic, thereby defining the rejection regions for the corresponding hypothesis tests. For example, when using the LRCLASS statistic to assess the parent-child relationship test at a FPR of $1.7\times 10^{-5}$, the suggested rejection region consists of values equal to or exceeding 6860.
Table 3: Parent-child (PC) and full-sibling (SB) test thresholds
Test statistic PC threshold at $\alpha=1.7\times 10^{-5}$ SB threshold at $\alpha=1.2\times 10^{-5}$
LRLAF 8251 10280
LRAVG 25111 28429
LRMAX 46070 79538
LRMIN 5666 5835
LRCLASS 6860 5835
[…]
Subsequently, Table 4 shows the power of each test for the parent-child relationship at an FPR of $1.7\times 10^{-5}$ and for the full-sibling relationship at an FPR of $1.2\times 10^{-5}$. Additionally, the table includes 95% confidence intervals for the power estimates to provide a more robust measure of precision.
Table 4: Power for the parent-child (PC) test at an FPR of $1.7\times 10^{-5}$ and for the full-sibling (SB) test at an FPR of $1.2\times 10^{-5}$
Statistic PC Power 95% CI SB Power 95% CI
LRLAF 0.8383 (0.83754, 0.83898) 0.5431 (0.54215, 0.54410)
LRAVG 0.7773 (0.77647, 0.77810) 0.5110 (0.51003, 0.51199)
LRMAX 0.7814 (0.78062, 0.78224) 0.4885 (0.48751, 0.48947)
LRMIN 0.8176 (0.81688, 0.81840) 0.5318 (0.53085, 0.53280)
LRCLASS 0.8116 (0.81079, 0.81233) 0.5748 (0.57388, 0.57582)
[…]
For the parent-child relationship test, as presented in Table 4, the LRLAF statistic demonstrates the highest power, achieving approximately 83.8%. This is followed by LRMIN and LRCLASS, both exhibiting power values in the range of 81–82%. In contrast, LRMAX and LRAVG yield comparatively lower power, approximately 78%.
The primary objective of this study, however, is to assess the performance of the proposed statistic, LRCLASS. In the context of parent-child relationship testing, LRCLASS ranks third, with a power that is 2.7% lower than the highest-ranking LRLAF and 0.6% lower than the second-ranking LRMIN. In the full-sibling relationship test, LRCLASS exhibits superior performance, achieving the highest power among the five evaluated statistics at a FPR of $1.2\times 10^{-5}$. Notably, LRCLASS surpasses the second-ranking statistic, LRLAF, by a margin of approximately 3.2%. The ranking of the statistics for the full-sibling case, from highest to lowest performance, is as follows: LRCLASS, LRLAF, LRMIN, LRAVG, and LRMAX.
To further investigate the performance of LRCLASS, power comparisons were visualized by plotting its power against each of the existing methods. The analysis was extended to include a range of FPRs, spanning from $10^{-6}$ to $4\times 10^{-5}$. From Figure 1, the results indicate that LRCLASS consistently outperforms LRMAX across nearly the entire range of FPRs considered. However, its performance relative to LRLAF, LRAVG, and LRMIN is less conclusive, as no clear dominance is observed throughout the FPR range analyzed. For the full-sibling relationship test, as shown in Figure 2, LRCLASS demonstrates superior performance compared to all other statistics across the entire range of FPRs considered.
Figure 1: Power comparison of LRCLASS to existing methods for parent-child testing
Figure 2: Power comparison of LRCLASS to existing methods for full-sibling testing
3.2 Classification Comparison
Tables 5 and 6 present the performance evaluation of each classification method using the validation dataset and the test dataset, respectively. The best-performing method for each criterion is emphasized in bold. It is important to note that the criteria listed in the first column of Tables 5 and 6 represent technical terms, which are explained here:
Table 5: Classification performance of each method on the validation dataset
Logistic Mthd A Logistic Mthd B Naive Bayes
Precision 0.6975 0.6919 0.6916
Recall 0.6985 0.6944 0.6938
Specificity 0.8500 0.8488 0.8511
Accuracy 0.8259 0.8237 0.8232
Error 0.1741 0.1763 0.1768
F1-score 0.6864 0.6824 0.6857
[…]
Table 6: Classification performance of each method on the test dataset
Logistic Mthd A Logistic Mthd B Naive Bayes
Precision 0.5003 0.4928 0.4924
Recall 0.4622 0.4584 0.4693
Specificity 0.7915 0.7896 0.7869
Accuracy 0.7219 0.7197 0.7237
Error 0.2781 0.2803 0.2763
F1-score 0.4435 0.4378 0.4462
[…]
First, we recall four standard terms: True Positive (TP) refers to a sample correctly identified as positive, while True Negative (TN) is a sample correctly identified as negative. False Positive (FP) occurs when a sample is incorrectly classified as positive, and False Negative (FN) occurs when a sample is incorrectly classified as negative. The metrics for classification are defined as follows:
  • Precision: the number of samples correctly classified as positive out of all samples predicted to be positive by the model. Mathematically, it is TP/(TP+FP). Since there are four subpopulations, each is treated as the positive class in turn, and precision is calculated for each. The average precision across the four subpopulations is then used as the final measure. This averaging approach is applied consistently to all other measures.
  • Recall: the number of samples correctly predicted to belong to the positive class out of all samples that actually belong to the positive class, written as TP/(TP+FN).
  • Specificity: the number of samples correctly predicted to be in the negative class out of all samples that actually belong to the negative class, given by TN/(FP+TN).
  • Accuracy: the number of samples correctly classified out of all samples in the test set, calculated as (TP+TN)/(TP+FP+TN+FN).
  • Error: the number of samples incorrectly classified out of all samples in the test set, given by (FP+FN)/(TP+FP+TN+FN).
[…]
As presented in Table 5, logistic regression with Method A demonstrates the best classification performance on the validation set across all evaluated criteria, except for specificity. Conversely, when classification is applied to the test dataset, as shown in Table 6, the Naive Bayes classifier outperforms in terms of recall, accuracy, error, and F1-score, while logistic regression with Method A performs best when evaluated using precision and specificity. It is also observed that the performance of logistic regression with Method B is slightly lower than that of Method A.
Lastly, Table 7 presents the confusion matrices for the three classification methods applied to the test dataset, where predicted class labels are compared against the actual (true) labels. The models categorize data points into one of four subpopulations: NO, NE, CT, or SO. Each diagonal entry in these matrices represents a correct classification, where the true and predicted subpopulations align.
Table 7: Confusion matrices for the three classification models on the test dataset
Logistic Mthd A Logistic Mthd B Naive Bayes
Truth Truth Truth
NO NE CT SO NO NE CT SO NO NE CT SO
Pred. NO 54 9 19 9 50 10 20 11 59 12 23 12
Pred. NE 84 219 75 84 87 218 75 84 90 239 90 96
Pred. CT 54 65 99 61 54 65 101 59 40 44 78 43
Pred. SO 9 10 19 56 10 10 16 55 13 9 20 59
Total 201 303 212 210 201 303 212 209 202 304 211 210
[…]
The overall classification accuracy is calculated by dividing the sum of the diagonal entries by the total number of observations in each confusion matrix. For logistic regression using Method A, the estimated correct classification rate is $(54+219+99+56)/926=46.22\%$. For logistic regression using Method B, the estimated rate is $(50+218+101+55)/925=45.84\%$. Finally, the estimated correct classification rate for Naive Bayes classification is $(59+239+78+59)/927=46.93\%$. Note that the original test dataset contained a total of 929 observations. However, due to technical constraints, some observations were excluded for specific methods. For logistic regression, exclusions occurred because certain alleles in the test dataset had never been observed in the training dataset, rendering the method inapplicable. In the case of Naive Bayes, the constraint arose from misalignment between the allele frequency data and the raw data. Specifically, while the allele frequency dataset included the probability of allele 34 at locus D21S11, the test dataset recorded this information as 34.1 or 34.2, creating ambiguity. As a result, such uncertain cases were excluded from the analysis.
4 Discussions
In terms of power comparison, as shown in Table 4, the performance can be categorized into two distinct groups. The first group includes LRLAF, LRMIN, and LRCLASS, all of which demonstrate strong performance in both the parent-child and full-sibling tests. The second group consists of LRAVG and LRMAX, which clearly exhibit lower power compared to the other statistics. This observation is consistent with the results of previous studies conducted on the Colorado population (Kooakachai et al. 2019). In particular, we recommend using the proposed statistic LRCLASS for the full-sibling test in the Thai population, as it consistently outperforms all other statistics, as illustrated in Figure 2. There are no significant limitations associated with using these test statistics, as the computational time required for each method is comparable.
From Tables 5 and 6, it is evident that the three classification methods demonstrate comparable performance in classifying DNA profiles. As presented in Table 6, the Naive Bayes classifier demonstrates superior performance across recall, accuracy, error, and F1-score, underscoring its suitability for tasks demanding well-rounded effectiveness across multiple metrics. In contrast, logistic regression with Method A excels in precision and specificity, making it particularly advantageous in contexts where minimizing false positive classifications is critical, such as scenarios where the cost or risk of false positives outweighs that of false negatives. Logistic regression with Method B exhibits consistent performance but slightly underperforms compared to both Naive Bayes and logistic regression with Method A across most evaluation metrics.
When evaluating the correct classification rate, the Naive Bayes classifier achieves the highest estimated rate at 46.93%, with a 95% confidence interval of (0.437, 0.502). However, this difference is not statistically significant when compared to the other methods. Logistic regression using Method A demonstrates an estimated correct classification rate of 46.22%, accompanied by a 95% confidence interval of (0.430, 0.495), which overlaps substantially with the performance of the Naive Bayes classifier. Overall, these findings suggest that the three classification methods exhibit comparable performance, with no single method showing a distinct advantage over the others.
As presented in Table 7, all classification methods achieve the highest accuracy in identifying DNA profiles from the NE class. Among them, the Naive Bayes classifier performs the best, correctly identifying 239 out of 304 NE samples, equating to approximately 78.6%. In contrast, the models struggle significantly with other classes, where the accuracy of correct classifications remains below 50% across all methods. This suggests that for individuals from the Thai population, DNA profiles from the northeastern region are the most consistently classified using both logistic regression and the Naive Bayes method, whereas profiles from other regions show considerably lower classification accuracy.
To discuss the pros and cons of each method, we first note that Naive Bayes is advantageous due to its simplicity and faster computation, as it does not require a training dataset, unlike logistic regression, which necessitates a training phase before classification can be applied. The Naive Bayes method benefits from reduced computational time; however, logistic regression may be preferable when working with specific subpopulations. For instance, as shown in Table 7, logistic regression performs better in classifying DNA profiles from the central region of Thailand. Specifically, Methods A and B of logistic regression correctly classified 99 and 101 samples, respectively, out of 212, achieving an accuracy of approximately 46–47%. In contrast, Naive Bayes yielded a lower accuracy of 36%, correctly classifying only 78 out of 211 samples. This discrepancy highlights the bias of the Naive Bayes method, which performs well for the Thai Northeastern subgroup but not for the Central subgroup.
Finally, among the two forms of logistic regression proposed, we recommend Method A, which uses 30 predictors by treating each allele across 15 loci as a separate predictor. This approach demonstrates superior overall performance, as evidenced by the results presented in Tables 5 and 6.
5 Conclusions
In this study, we introduced a novel approach to familial DNA testing, specifically designed for evaluating parent-child and full-sibling relationships with a population substructure. The proposed method LRCLASS employs a two-step process: classification is performed first, and the resulting classifications are subsequently used to compute the likelihood ratio. Our findings highlight the exceptional performance of the LRCLASS statistic, particularly when paired with the Naive Bayes classifier, which demonstrates superior power in full-sibling relationship testing within the Thai population compared to all existing statistics. Additionally, logistic regression methods were investigated as alternative classification tools, with their performance shown to be comparable to that of the Naive Bayes classifier.
We propose several potential avenues for future research to build upon the findings of this study. First, while the LRCLASS, LRLAF, and LRMIN statistics demonstrate strong performance, it remains an open question whether they represent the most powerful approaches for relatedness testing. Identifying a new statistic or even more complex tools for classification that consistently outperform existing ones could be transformative. Since there is currently no theoretical framework for comparing the efficacy of these statistics, further methodological advancements and theoretical developments in this area would be invaluable.
Second, although our hypothesis tests primarily focused on parent-child and full-sibling relationships, future research could expand to include more complex familial relationships, such as half-siblings, first cousins, second cousins, and double first cousins. Examining how statistical power varies with the complexity of the relationship could provide deeper insights into the capabilities and limitations of current methods.
Finally, our analysis was conducted under the assumption of the Thai population with its distinct regional substructure. Extending this analysis to other populations with varying substructures would be a meaningful step, enabling cross-population comparisons and shedding light on how population-specific genetic diversity impacts the performance of relatedness testing statistics. Such work could contribute to a broader understanding of the generalizability and adaptability of these methods across diverse genetic landscapes.
6 List of abbreviations
DNA Deoxyribonucleic Acid
LR Likelihood ratio
LRLAF Likelihood ratio statistic involving the use of average allele frequency distribution
LRMAX Likelihood ratio statistic based on the maximum of all available likelihood ratios
LRMIN Likelihood ratio statistic based on the minimum of all available likelihood ratios
LRAVG Likelihood ratio statistic based on the average of all available likelihood ratios
LRCLASS Likelihood ratio statistic involving a classification step
Y-STR Short tandem repeat on the Y-chromosome
STR Short tandem repeat
USA United States of America
A Adenine
T Thymine
G Guanine
C Cytosine
HVR1 Hypervariable region 1
CIFS Central Institute of Forensic Science
NO North
NE Northeastern
CT Central
SO South
IBD Identical by descent
Mb Megabase
prob Probability
subpop Subpopulation
FPR False positive rate
PC Parent-child
SB Full-sibling
CI Confidence interval
TP True positive
TN True negative
FP False positive
FN False negative
Mthd Method
Pred Predict
7 Declarations
Ethics approval and consent to participate
The dataset used in this study was approved by the ethics committees of the Faculty of Medicine, Ramathibodi Hospital, Mahidol University, Thailand (COA No. MURA2023/899). The requirement for informed consent was waived for each participant.
Consent for publication
Not applicable
Availability of data and material
The training datasets generated during the current study, along with the test dataset, are available from the corresponding author upon reasonable request. The allele frequency distributions, which serve as the basis for the generated datasets, can be found in the public literature (Shotivaranon et al. 2009).
Competing interests
The authors declare that they have no competing interests.
Funding
This research did not receive any financial support from institutions or organizations.
Authors’ contributions
AJ Conceptualization, methodology, software, formal analysis for the classification comparison part, investigation, writing – original draft. CL Conceptualization, methodology, software, formal analysis for the power comparison part, investigation. BR Conceptualization, methodology, dataset provision. JS Conceptualization, methodology, dataset provision. MK Conceptualization, methodology, software, writing – review and editing, supervision, project administration. All authors read and approved the final manuscript.
Acknowledgments
The authors wish to express their gratitude to the editor, associate editor, and referees for their valuable comments and suggestions, which significantly enhanced the quality of this article.